Sue consumes only two goods, food and clothing. The marginal utility on food is 22, and the marginal utility on clothing is 9. The price of food is $2.00/unit, and the price of clothing is $1.00/unit. Is Sue maximizing her utility? If not, she is consuming too much of food or clothing?
Answer #1

Answer

Utility maximizing condition:

In order to maximize utility a consumer should consume that quantity at which MUx/Px = MUy/Py, where MU represents Marginal Utility and P represents price.

Hence Here MUf/Pf = 22/2 = 11 and MUc/Pc = 9/1 = 9.

Hence MUf/Pf is not equal to MUc/Pc

Hence Currently Sue is not maximizing her utility.

MUf/Pf > MUc/Pc

According to Law of diminishing marginal utility Marginal utility derive from additional utility must decline.

i.e. As consumption increases Marginal Utility decreases..

As, MUf/Pf > MUc/Pc , Hence MUc should be more and hence she should decrease the consumption of Clothing. Hence Currently she is consuming too much of Clothing